# OpenEvents V1: Large-Scale Benchmark Dataset for Multimodal Event Grounding

### v1

@@ Abstract.
We introduce OpenEvents V1, a large-scale benchmark dataset aimed at advancing event-centric vision-language understanding. Unlike conventional image captioning and retrieval datasets that emphasize surface-level descriptions, OpenEvents V1 focuses on contextual and temporal grounding through two primary tasks: (1) generating rich, event-aware image captions and (2) retrieving event-relevant images based on narrative-style textual queries. The dataset contains over 200,000 news articles and 400,000 associated images sourced from CNN and The Guardian, spanning diverse domains and time periods. We provide extensive baseline results and standardized evaluation protocols for both tasks. OpenEvents V1 establishes a robust foundation for developing multimodal models capable of deep reasoning over complex real-world events. The dataset is available at https://ltnghia.github.io/eventa/openevents-v1.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrievalccs: Applied computing Media artsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning
@@
((b)) Event-based image retrieval
Figure 1. OpenEvents V1 dataset supports two tasks for event grounding.
@@
Two subfigures: (a) … (b) …
1. Introduction
Figure 2. Sample image-caption pairs in our OpenEvents V1 dataset.
@@ 1. Introduction
Problem 1 – Event-enriched image captioning. Given an input image, models are required to retrieve relevant contextual information from an external article database and generate a detailed, context-aware caption. This task encourages models to go beyond surface-level visual cues and incorporate real-world knowledge such as named entities, temporal context, and event outcomes. By grounding the caption in both visual content and textual evidence, the generated description reflects a more comprehensive understanding of the depicted event, supporting applications in visual journalism, archival, and automated media reporting.
Problem 2 – Event-based image retrieval. Given a caption describing a real-world event, participants must retrieve the corresponding image from a large-scale database. Unlike traditional keyword-matching systems, this task requires semantic alignment between narrative-style text and event-centric visual representations. It emphasizes the ability of models to reason across modalities and handle diverse linguistic expressions of the same event. This capability is crucial for information retrieval in domains such as news curation, multimedia search engines, and digital content recommendation.
@@ 4.1. Problem 1: Event-Enriched Image Captioning
4.2. Problem 2: Event-Based Image Retrieval
We evaluate the capability of models to retrieve the correct image from a large-scale visual database given a caption that narrates a real-world event. Unlike conventional retrieval tasks, where captions typically describe only the visual content, our setting presents a greater challenge: the caption is generated based on both the image and its associated article text. As a result, the descriptions often reflect deeper event semantics, such as named entities, temporal references, and contextual outcomes that may not be visually evident. This makes direct retrieval significantly harder, as models must align complex, narrative-style captions with visual content that is only partially represented in the text. Table 1 presents captioning performance across different models, with and without access to article text. The results indicate a clear improvement when models are provided with additional contextual information from the associated article.
Table 2. Problem 2 – Retrieval results on the public test set.
@@ 4.2. Problem 2: Event-Based Image Retrieval
We evaluate retrieval performance using standard ranking metrics: Nearest Neighbor Accuracy (NN) (Cover and Hart, 1967), Normalized Discounted Cumulative Gain (NDCG) (Wang et al., 2013), mean Average Precision (mAP), and Area Under the Precision-Recall Curve (AUC) (Davis and Goadrich, 2006).
@@
5. Applications
The OpenEvents V1 benchmark is designed to support a broad range of real-world and research-driven applications. Its structure, scale, and dual-task focus make it particularly suitable for evaluating and advancing multimodal understanding in complex, event-rich environments.
Event-Enriched Image Captioning.
Our dataset enables the development of captioning systems that move beyond object-level description to generate narrative-style summaries that incorporate temporal, causal, and contextual information. This has strong implications for media analysis, automated journalism, and knowledge grounding.
Event-Based Image Retrieval.
By supporting realistic, information-rich queries and large-scale search space, OpenEvents V1 facilitates the training and evaluation of retrieval models that bridge the gap between narrative text and real-world visual content.
Retrieval-Augmented Generation
The benchmark provides both image-caption pairs and a structured article database, making it an ideal testbed for retrieval-augmented captioning systems, including RAG, BLIP-Retrieval, and hybrid LLM architectures.
Multimodal Reasoning and QA.
The question-answering intermediate stage can be adapted for visual QA, entity extraction, and reasoning tasks, especially where grounding in external factual context is essential.
Educational and Journalistic Tools.
The dataset also enables the development of applications such as automatic article summarizers, timeline generation tools, and visual story reconstruction engines.
@@ 6. Conclusion
We introduced OpenEvents V1, a large-scale benchmark for event-enriched image captioning and event-based image retrieval. The dataset consists of over 400,000 images and 200,000 articles collected from reputable news sources across a 15-year span. To generate high-quality captions, we propose a five-stage human-agentic framework that integrates vision-language models with article grounding and final human verification. In addition, we introduce two benchmark tasks: Event-Enriched Image Captioning and Event-Based Image Retrieval. We also present baseline results for both tasks. This benchmark fills a critical gap in multimodal research by providing large-scale, real-world supervision enriched with contextual information.
We expect that OpenEvents V1 will serve as a foundation for future research in grounded generation, retrieval, event understanding, and cross-modal reasoning. Our benchmark and baselines will be made publicly available to facilitate reproducibility, comparison, and future challenge organization.

### v2

@@ Abstract.
We introduce OpenEvents V1, a large-scale benchmark dataset designed to advance event-centric vision–language understanding. Unlike conventional image captioning and retrieval datasets that focus on surface-level descriptions, OpenEvents V1 dataset emphasizes contextual and temporal grounding through three primary tasks: (1) generating rich, event-aware image captions, (2) retrieving event-relevant news articles from image queries, and (3) retrieving event-relevant images from narrative-style textual queries. The dataset comprises over 200,000 news articles and 400,000 associated images sourced from CNN and The Guardian, spanning diverse domains and time periods. We provide extensive baseline results and standardized evaluation protocols for all tasks. OpenEvents V1 establishes a robust foundation for developing multimodal AI systems capable of deep reasoning over complex real-world events. The dataset is publicly available at https://ltnghia.github.io/eventa/openevents-v1.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3758264ccs: Information systems Information retrievalccs: Applied computing Media artsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning
Figure 1. Sample image-caption pairs in the proposed OpenEvents V1 dataset.
1. Introduction
@@
((b)) Event-based article retrieval
((c)) Event-based image retrieval
Figure 2. OpenEvents V1 dataset supports three tasks for event grounding.
@@
Problem 1 – Event-enriched image captioning. Given an input image, models are required to retrieve relevant contextual information from an external article database and generate a detailed, context-aware caption (2(a)). This task encourages models to go beyond surface-level visual cues and incorporate real-world knowledge such as named entities, temporal context, and event outcomes. By grounding the caption in both visual content and textual evidence, the generated description reflects a more comprehensive understanding of the depicted event, supporting applications in visual journalism, archival, and automated media reporting.
Problem 2 – Event-based article retrieval. In practical deployments, an image captioning system may be provided with only the image, lacking the corresponding article. Because event-enriched image captioning requires both visual cues and contextual knowledge, the first step is to recover that missing context. To address this, we incorporate an article retrieval stage that locates the most relevant news article for the given image, ensuring the captioning model has access to the background information needed to produce accurate, event-aware descriptions (2(b)).
Problem 3 – Event-based image retrieval. Given a caption describing a real-world event, participants must retrieve the corresponding image from a large-scale database (2(c)). Unlike traditional keyword-matching systems, this task requires semantic alignment between narrative-style text and event-centric visual representations. It emphasizes the ability of models to reason across modalities and handle diverse linguistic expressions of the same event. This capability is crucial for information retrieval in domains such as news curation, multimedia search engines, and digital content recommendation.
@@ 4.1. Problem 1: Event-Enriched Image Captioning
4.2. Problem 2: Event-Based Article Retrieval
Table 2. Problem 2 – Article retrieval results on the public test set.
Method mAP NDCG NN AUC
CLIP 0.3504 0.4594 0.23 0.0575
OPEN CLIP 0.3141 0.4249 0.2016 0.0487
We evaluate CLIP(Radford et al., 2021) and OpenCLIP(Cherti et al., 2023) on four ranking-based metrics, such as mean Average Precision (mAP), Normalized Discounted Cumulative Gain (NDCG), Nearest Neighbor accuracy (NN), and Area Under the Precision–Recall Curve (AUC). Results in Table 2 show a consistent advantage for CLIP across all measures, suggesting superior alignment with the event-centric properties of OpenEvents V1.
This performance difference may be attributed to differences in training data. CLIP, the model released by OpenAI, was trained on a carefully curated dataset with strong image–text correspondence. OpenCLIP, while architecturally similar, was trained on larger but less targeted public datasets, such as LAION (Schuhmann et al., 2022), DataComp (Gadre et al., 2023), CommonPool (Ostrom, 2002), and YFCC100M (Thomee et al., 2016), that may not reflect the style and content of journalistic imagery. This difference in data curation likely explains CLIP’s stronger performance, reinforcing the broader insight that domain alignment in training data is critical for high retrieval accuracy in event-focused tasks.
4.3. Problem 3: Event-Based Image Retrieval
We evaluate the capability of models to retrieve the correct image from a large-scale visual database given a caption that narrates a real-world event. Unlike conventional retrieval tasks, where captions typically describe only the visual content, our setting presents a greater challenge: the caption is generated based on both the image and its associated article text. As a result, the descriptions often reflect deeper event semantics, such as named entities, temporal references, and contextual outcomes that may not be visually evident. This makes direct retrieval significantly harder, as models must align complex, narrative-style captions with visual content that is only partially represented in the text.
Table 3. Problem 3 – Image retrieval results on the public test set.
@@ 5. Conclusion
In this paper, we introduced OpenEvents V1, a large-scale benchmark designed to advance research in event-enriched image captioning and event-based image retrieval. To ensure high-quality outputs for event-enriched image captioning, we proposed a five-stage human–agentic framework that integrates vision–language models with article grounding and final human verification. This framework supports three benchmark tasks: event-enriched image captioning, event-based article retrieval, and event-based image retrieval. Baseline results for all tasks demonstrate the challenges and opportunities presented by this dataset, filling a critical gap in multimodal research through large-scale, real-world supervision enriched with contextual information. Our benchmark and baselines have been made publicly available to facilitate reproducibility and comparison. This dataset also has been utilized in the ACM Multimedia Grand Challenge on Event-Enriched Image Analysis (EVENTA’25)(Tran et al., 2025) to encourage the development of context-aware multimodal AI systems.
The OpenEvents V1 dataset can support a wide spectrum of real-world and research-oriented applications. Its dual focus on captioning and retrieval makes it an ideal testbed for evaluating retrieval-augmented generation systems and hybrid LLM architectures. The integrated article database provides a robust foundation for applications where external factual grounding is essential, enabling adaptations for visual question answering, entity extraction, and multimodal reasoning.
Future work will focus on expanding the benchmark with multilingual content, multimodal event graphs, and video-based event understanding to capture richer temporal dynamics.